# Disentangling Reasoning and Knowledge in Medical Large Language Models

### v1

@@ Abstract
Medical reasoning in large language models (LLMs) aims to replicate clinicians’ cognitive processes when interpreting patient data and making diagnostic decisions. However, evaluating true reasoning capabilities remains challenging, as widely used benchmarks-such as MedQA-USMLE, MedMCQA, and PubMedQA-often conflate questions requiring medical reasoning with those solvable through factual recall. We address this limitation by systematically disentangling reasoning-heavy from knowledge-heavy questions across 11 biomedical QA benchmarks using a PubMedBERT-based classifier that achieves human-level performance (81%). Our analysis reveals that only 32.8% of benchmark questions involve complex reasoning, with the majority focused on factual understanding. Using this stratified dataset, we evaluate recent biomedical reasoning models (HuatuoGPT-o1, MedReason, m1) alongside general-domain models (DeepSeek-R1, o4-mini, Qwen3) and observe a consistent performance gap between knowledge and reasoning—for example, m1 scores 60.5% vs. 47.1%, respectively. To assess robustness, we conduct adversarial evaluations where models are prefilled with incorrect answers before being asked to reconsider. Biomedical models show substantial degradation in this setting (e.g., MedReason drops from 44.4% to 29.3%), while RL-trained and larger general-domain models are more resilient. Based on these insights, we train BioMed-R1-8B using supervised fine-tuning and reinforcement learning on reasoning-heavy examples. While it achieves the strongest overall and adversarial performance among similarly sized models, there remains ample room for improvement. Incorporating additional reasoning-rich data sources, such as clinical case reports, and training on adversarial or backtracking scenarios—with reinforcement learning to encourage self-correction—may further enhance robustness and reliability [^1] Codebase available at https://github.com/togethercomputer/BioMed-R1.
@@ 1 Introduction
In this paper, we disentangle 11 biomedical evaluation benchmarks into reasoning-heavy and knowledge-heavy subsets by fine-tuning a PubMedBERT classifier to label each question. The classifier achieves human-level performance (~81%) when validated against both gold-standard annotations and independent clinical experts. Our analysis reveals that only 32.8% of the benchmark questions considered are reasoning-heavy. Using this annotated dataset, we evaluate recent biomedical reasoning models (MedReason, m1, HuatuoGPT-o1), consistently finding a performance gap between knowledge-heavy and reasoning-heavy questions—for example, m1 scores 60.5% vs. 47.1%, respectively. While factual knowledge can often be grounded through external tools like retrieval-augmented generation (RAG) or verified via resources such as UpToDate or Google Search, clinical reasoning lacks such external anchors. It requires models to synthesize nonspecific symptoms, weigh competing diagnoses, and make contextual decisions—capabilities that cannot be easily offloaded to retrieval systems. As such, reasoning represents a more fundamental and high-stakes benchmark for biomedical LLMs. Finally, in a robustness testing setup where models are prefixed with incorrect answers, biomedical models degrade substantially (e.g., MedReason drops from 44.4% to 29.3%), while RL-trained general-domain models demonstrate greater resilience.
Based on these insights, we identify the m23k dataset as the strongest publicly available source of medical reasoning supervision and show that augmenting it with RL on 7,627 high-reasoning examples—filtered from MedQA-USMLE, MedMCQA, and HeadQA [33] using our BERT classifier—improves m1’s reasoning accuracy from 47.1% to 49.1%. We refer to this model as BioMed-R1-8B, which achieves the best overall and adversarial performance among all models of similar scale. While our findings offer early evidence that RL on reasoning-heavy examples can enhance both reasoning and knowledge accuracy—as well as adversarial robustness—there remains substantial room for improvement. For instance, incorporating additional training data from other reasoning-rich sources, such as clinical case reports, could further boost performance. Likewise, training models on carefully constructed adversarial traces and backtracking scenarios, reinforced through RL, may further improve robustness and reliability.
@@ 3.1 Evaluation Data Curation
Representative examples of knowledge- and reasoning-heavy questions are shown in Figure 1. Reasoning-heavy questions tend to be much longer than knowledge-heavy ones. We illustrate this difference through the word count distribution in Supplementary Figure S4. To test this correlation, we trained a logistic regression model that uses word count as input to predict reasoning probability. As shown in Figure 3, this simple baseline achieves a notable 72% accuracy. While the PubMedBERT classifier significantly outperforms it, the baseline supports the observation that reasoning questions generally require longer problem statements—likely because they demand synthesis of more information. We further present representative correct and failure cases of our classifier in Supplementary Figures S9, S10 and S11.
@@ 4.1 Model Training
While recent medical reasoning models such as HuatuoGPT-o1 [6], MedReason [37], and m1 [13] publicly release their fine-tuned models, we chose to retrain all models from scratch using their respective SFT datasets. This decision was driven by three main considerations. First, some released models include additional RL steps beyond SFT (e.g., HuatuoGPT-o1), complicating comparisons of reasoning capabilities learned mostly through SFT. Second, several training datasets contain benchmark test questions—such as MedReason from MedXpert and HLE, or m1 from PubMedQA—raising concerns about data leakage. Third, retraining all models on a unified, decontaminated SFT corpus under a standardized training pipeline enables fair and reproducible comparisons between architectures, including Llama-3.1-8B-Instruct [9] and Qwen2.5-7B-Instruct [38]. To support this, we filtered the SFT datasets to exclude any known overlaps with evaluation benchmarks and trained all models using the same framework. Full dataset and training details are provided in Appendices B and A.
@@ 4.3 Main Evaluations
Normal Adv. (diff) Normal Adv. (diff) Normal Adv. (diff)
@@
How Do General-Domain Reasoning Models Perform on Medical Benchmarks? Before evaluating biomedical reasoning models, we first assess the performance of leading open-source general-domain reasoning models—including Qwen3-235B, DeepSeek-R1, GPT-4o, and o4-mini—on our benchmark setup. This is important because many medical reasoning datasets, such as m23k, medical-o1, and MedReason, are distilled from these models, making their capabilities directly relevant to the foundations of biomedical instruction tuning. As shown in Table 1, general-domain models consistently perform better on knowledge-heavy questions than on reasoning-heavy ones. However, while reasoning accuracy still lags behind knowledge accuracy, these general-domain models achieve substantially stronger reasoning performance overall and exhibit greater robustness under adversarial conditions. Notably, larger RL-trained models like DeepSeek-R1 and o4-mini show better robustness even under adversarial prompts—where the model is prefixed with a misleading hypothesis. These findings highlight both the current ceiling for open-domain reasoning performance and the potential for improving biomedical models through cross-domain transfer or stronger supervision signals.
@@
Table 2: Accuracy across knowledge, reasoning, and overall categories for Llama-3.1-8B-Instruct variants on 11 biomedical QA benchmarks. Subscripts denote 95% confidence intervals. Reasoning accuracy is consistently lower than knowledge. Notably, CoT prompting, retrieval-based methods, and SFT without reasoning traces remain competitive with fully fine-tuned models. See Supplementary Figure S6 for the CoT prompt and Supplementary Table S3 for task-level scores.
@@
RAG methods also yield strong results. The best-performing variant, which retrieves from MedReason, achieves 43.5% on reasoning and 52.8% overall—surpassing simple CoT prompting on both metrics. Notably, the RAG baseline on Medical o1 performs nearly as well as its fully SFT counterpart (51.1% vs. 51.9%) and even outperforms it on MedReason (52.8% vs. 50.4%). These results suggest that access to retrieved reasoning traces, even without gradient updates, can help the model reason more effectively at inference time by providing structured exemplars to guide decision-making.
@@
Together, these results suggest that while fine-tuning on high-quality reasoning traces yields the best performance, carefully chosen test-time strategies—especially CoT and RAG—can substantially narrow the gap, often matching or exceeding models trained on SFT datasets. A task-wise breakdown of performance is provided in Supplementary Tables S3 and S4.
@@
Llama w/ m23k SFT 60.559.5 - ${}_{\text{59.5 - 61.5}}$ 47.145.7 - ${}_{\text{45.7 - 48.7}}$ 56.155.3 - ${}_{\text{55.3 - 56.9}}$
+ Knowledge Hard SFT 60.159.1 - ${}_{\text{59.1 - 61.2}}$ 47.746.1 - ${}_{\text{46.1 - 49.3}}$ 56.055.1 - ${}_{\text{55.1 - 56.9}}$
+ Reason Hard SFT 60.559.5 - ${}_{\text{59.5 - 61.6}}$ 49.147.6 - ${}_{\text{47.6 - 50.7}}$ 56.855.9 - ${}_{\text{55.9 - 57.7}}$
+ Reason Hard RL (BioMed-R1-8B) 63.662.5 - ${}_{\text{62.5 - 64.7}}$ 49.047.5 - ${}_{\text{47.5 - 50.6}}$ 58.857.9 - ${}_{\text{57.9 - 59.6}}$
Table 3: Performance of Llama-3.1-8B-Instruct variants fine-tuned on m23k, then further trained with either SFT or RL (GRPO) on 7,627 reasoning-hard examples. A control using knowledge-hard examples is also included. Adding reasoning-hard data improves reasoning accuracy more than knowledge-hard data, while RL further boosts both reasoning and knowledge performance.
Including Hard Reasoning Problems in Training Mixture Enhances Reasoning. Building on our finding that Llama fine-tuned on the m23K dataset yields the strongest performance, we next investigated whether carefully curated reasoning examples could further enhance the model’s performance. We combined three commonly used training datasets—MedQA, MedMCQA, and HeadQA—and selected 7,627 “reasoning-hard” multiple-choice questions that met the following criteria: (i) neither Llama-3.1-8B-Instruct nor Qwen-2.5-7B-Instruct answered them correctly more than 50% of the time, (ii) our PubMedBert classifier assigned a reasoning probability above 0.75, and (iii) they did not overlap with the m23K dataset. For comparison, we also curated a set of 7,627 “knowledge-hard” questions that met the same difficulty threshold but were classified as knowledge-heavy (reasoning probability below 0.75). We generated detailed reasoning traces for both sets using R1 and trained two separate models: one mixing m23K with the reasoning-hard set (“+ Reason Hard SFT”) and another mixing m23K with the knowledge-hard set (“+ Knowledge Hard SFT”). As shown in Table 3, incorporating reasoning-hard examples led to a substantial improvement in reasoning accuracy (47.1% → 49.1%), while adding knowledge-hard examples resulted in negligible gains.
Finally, we applied the 7,627 reasoning-hard examples in a reinforcement learning (RL) setup using GRPO (“+ Reason Hard RL”), continuing training from the m23K SFT checkpoint. RL training was conducted using the EasyR1 framework [41, 29], with key hyperparameters detailed in Supplementary Table S1. The resulting model, referred to as BioMed-R1-8B (RL), achieved 58.8% overall accuracy and 63.6% on knowledge questions, demonstrating that policy optimization on reasoning-hard examples provides greater gains than SFT alone.
Two interesting observations emerge. First, while SFT on reasoning-hard examples mainly improves reasoning accuracy, RL boosts performance on both reasoning- and knowledge-heavy tasks. Second, as shown in Supplementary Table S7, BioMed-R1-8B (RL) generalizes well to out-of-distribution knowledge-heavy questions, whereas the reasoning improvements appear mostly in the in-domain setting. Although the exact reason for RL’s superior performance on knowledge-heavy tasks is unclear, one plausible explanation is that reasoning-heavy questions still rely on accurate factual recall. SFT, heavily conditioned on verbose R1 reasoning traces, may constrain the model’s ability to optimize for concise factual correctness. In contrast, RL—by directly optimizing for final answer accuracy—may encourage more effective integration of both reasoning and factual knowledge. While preliminary, these results suggest that emphasizing difficult reasoning cases—and applying RL—can meaningfully enhance both reasoning and knowledge capabilities of biomedical LLMs. A detailed benchmark-wise breakdown is provided in Supplementary Table S6.
Model Knowledge Reasoning Overall
Normal Adv. (diff) Normal Adv. (diff) Normal Adv. (diff)
@@
Interestingly, models trained with RL—specifically HuatuoGPT-o1-8B (RL) and BioMed-R1 (RL)—exhibit the least performance degradation compared to both standard SFT models and their own SFT counterparts (HuatuoGPT-o1-8B (SFT), BioMed-R1 (SFT)). These results suggest that RL may provide greater robustness to misleading initial reasoning and improve backtracking behavior. Another interesting observation is that, reasoning-heavy questions tend to suffer larger performance drops than knowledge-heavy ones, as evidenced by the higher relative percentage decreases across most models. This pattern further highlights a fundamental gap in current models’ reasoning capabilities: while they may perform well on fact-based queries, their ability to reason through uncertainty, reconsider initial impressions, and course-correct remains limited.

### v2

# Disentangling Knowledge and Reasoning in Medical Large Language Models

@@ Abstract
Medical reasoning in large language models aims to replicate clinicians’ cognitive processes when interpreting patient data and making diagnostic decisions. However, widely used benchmarks—such as MedQA-USMLE, MedMCQA, and PubMedQA—mix questions that require multi-step reasoning with those answerable through factual recall, complicating reasoning evaluation. To address this, we develop a PubMedBERT-based classifier (81% agreement with expert annotations) to disentangle reasoning-heavy from knowledge-heavy questions across 11 biomedical QA benchmarks, revealing that only 32.8% require complex reasoning. Using this stratification, we evaluate biomedical models (HuatuoGPT-o1, MedReason, m1) and general-domain models (DeepSeek-R1, o4-mini, Qwen3), and consistently observe lower performance on reasoning versus knowledge (e.g., HuatuoGPT-o1: 56.9% vs. 44.8%). To assess robustness, we conduct adversarial evaluations where models are prefilled with incorrect answers before being asked to reconsider. Biomedical models show substantial degradation in this setting (e.g., MedReason drops from 50.4% to 24.4%), while RL-trained and larger general-domain models are more resilient. Performance declines more on reasoning-heavy questions, highlighting the brittleness of current medical reasoning capabilities. Based on these insights, we train BioMed-R1 models using supervised fine-tuning and reinforcement learning on reasoning-heavy and adversarial examples, encouraging self-correction and backtracking. Our models achieve the strongest overall and adversarial performance among similarly sized biomedical LLMs, yet ample room for improvement remains. Incorporating additional reasoning-rich data sources—such as clinical case reports—and developing training strategies that promote reasoning under uncertainty may further enhance robustness and diagnostic reliability [^1] Codebase available at https://github.com/zou-group/BioMed-R1.
@@ 1 Introduction
In this paper, we disentangle 11 biomedical evaluation benchmarks into reasoning-heavy and knowledge-heavy subsets by fine-tuning a PubMedBERT classifier to label each question. The classifier achieves human-level performance (~81%) when validated against both gold-standard annotations and independent clinical experts. Our analysis reveals that only 32.8% of the benchmark questions considered are reasoning-heavy. Using this annotated dataset, we evaluate recent biomedical reasoning models (MedReason, m1, HuatuoGPT-o1), consistently finding a performance gap between knowledge-heavy and reasoning-heavy questions—for example, HuatuoGPT-o1 scores 56.9% vs. 44.8%, respectively. While factual knowledge can often be grounded through external tools like retrieval-augmented generation (RAG) or verified via resources such as UpToDate or Google Search, clinical reasoning lacks such external anchors. It requires models to synthesize nonspecific symptoms, weigh competing diagnoses, and make contextual decisions—capabilities that cannot be easily offloaded to retrieval systems. As such, reasoning represents a more fundamental and high-stakes benchmark for biomedical LLMs.
To further probe model reliability, we introduce a robustness test where models are prefixed with incorrect answers and asked to reconsider. Biomedical models degrade substantially in this setting (e.g., MedReason drops from 50.4% to 24.4%), while RL-trained and larger general-domain models show greater resilience. The performance decline is more pronounced on reasoning-heavy questions, underscoring the brittleness of current medical reasoning capabilities. Encouragingly, smaller biomedical models can recover performance with minimal training on adversarial traces.
Motivated by these insights, we identify the m23k dataset as the strongest publicly available source of medical reasoning supervision. We augment it with RL on 7,627 high-reasoning questions—filtered from MedQA-USMLE, MedMCQA, and HeadQA [38] using our classifier—and incorporate a small number of adversarial traces during SFT. This improves both knowledge and reasoning accuracy, as well as adversarial robustness. Among training strategies, we find that combining SFT with RL consistently outperforms pure SFT or RL alone. Interestingly, RL enhances not only reasoning but also factual knowledge performance. We train two model variants: BioMed-R1-8B, based on Llama-3.1-8B-Instruct, and BioMed-R1-32B, based on Qwen2.5-32B-Instruct. BioMed-R1-8B achieves the best overall and adversarial performance among similarly sized models, while BioMed-R1-32B surpasses all other biomedical models in its class—including MedGemma-27B [12]—and rivals HuatuoGPT-o1-70B, despite being less than half its size and trained on significantly less data.
While our results show that RL on reasoning-heavy and adversarial samples improves model robustness and accuracy, there remains ample room for improvement. Future work may incorporate additional reasoning-rich data sources (e.g., clinical case reports), and further refine adversarial and backtracking-based training strategies. Such efforts could help build more reliable and self-correcting biomedical reasoning models.
@@ 2 Related Work
A recent study introduces an inference-time framework that separates memory recall from reasoning using control tokens, improving both performance and interpretability [21]. While their work focuses on disentangling these processes during generation, we argue that it is equally important to do so during evaluation—by assessing whether benchmarks truly measure reasoning or merely factual recall. Both efforts are complementary in advancing a more precise understanding of LLM reasoning capabilities.
@@ 3.1 Evaluation Data Curation
Representative examples of knowledge- and reasoning-heavy questions are shown in Figure 1. Reasoning-heavy questions tend to be much longer than knowledge-heavy ones. We illustrate this difference through the word count distribution in Supplementary Figure S4. To test this correlation, we trained a logistic regression model that uses word count as input to predict reasoning probability. As shown in Figure 3, this simple baseline achieves a notable 72% accuracy. While the PubMedBERT classifier significantly outperforms it, the baseline supports the observation that reasoning questions generally require longer problem statements—likely because they demand synthesis of more information. We further present representative correct and failure cases of our classifier in Supplementary Figures S10, S11 and S12.
@@ 4.1 Model Training
While recent medical reasoning models such as HuatuoGPT-o1 [7], MedReason [42], and m1 [17] publicly release their fine-tuned models, we chose to retrain all models from scratch using their respective SFT datasets. This decision was driven by three main considerations. First, some released models include additional RL steps beyond SFT (e.g., HuatuoGPT-o1), complicating comparisons of reasoning capabilities learned mostly through SFT. Second, several training datasets contain benchmark test questions—such as MedReason from MedXpert and HLE, or m1 from PubMedQA—raising concerns about data leakage. Third, retraining all models on a unified, decontaminated SFT corpus under a standardized training pipeline enables fair and reproducible comparisons between architectures, including Llama-3.1-8B-Instruct [13] and Qwen2.5-7B-Instruct [43]. To support this, we filtered the SFT datasets to exclude any known overlaps with evaluation benchmarks and trained all models using the same framework. Full dataset and training details are provided in Appendices A and B.
@@ 4.3 Main Evaluations
Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$)
@@
General-Domain Reasoning Models Remain Competitive on Medical Benchmarks. Before evaluating biomedical reasoning models, we first assess the performance of leading open-source general-domain reasoning models—including Qwen3-235B, DeepSeek-R1, GPT-4o, and o4-mini—on our benchmark setup. This is important because many medical reasoning datasets, such as m23k, medical-o1, and MedReason, are distilled from these models, making their capabilities directly relevant to the foundations of biomedical instruction tuning. As shown in Table 1, general-domain models consistently perform better on knowledge-heavy questions than on reasoning-heavy ones. However, while reasoning accuracy still lags behind knowledge accuracy, these general-domain models achieve substantially stronger reasoning performance overall and exhibit greater robustness under adversarial conditions. Larger RL-trained models like DeepSeek-R1 and o4-mini show better robustness even under adversarial prompts—where the model is prefixed with a misleading hypothesis. These findings highlight both the current ceiling for open-domain reasoning performance and the potential for improving biomedical models through cross-domain transfer or stronger supervision signals. A detailed task-level performance breakdown is provided in Supplementary Table S2.
@@
Table 2: Accuracy across knowledge, reasoning, and overall categories for Llama-3.1-8B-Instruct variants on 11 biomedical QA benchmarks. Subscripts denote 95% confidence intervals. Reasoning accuracy is consistently lower than knowledge. CoT prompting, retrieval-based methods, and SFT without reasoning traces remain competitive with fully fine-tuned models. See Supplementary Figure S7 for the CoT prompt and Supplementary Table S4 for task-level scores.
@@
RAG methods also yield strong results. The best-performing variant, which retrieves from MedReason, achieves 43.5% on reasoning and 52.8% overall—surpassing simple CoT prompting on both metrics. The RAG baseline on Medical o1 performs nearly as well as its fully SFT counterpart (51.1% vs. 51.9%) and even outperforms it on MedReason (52.8% vs. 50.4%). These results suggest that access to retrieved reasoning traces, even without gradient updates, can help the model reason more effectively at inference time by providing structured exemplars to guide decision-making.
@@
Together, these results suggest that while fine-tuning on high-quality reasoning traces yields the best performance, carefully chosen test-time strategies—especially CoT and RAG—can substantially narrow the gap, often matching or exceeding models trained on SFT datasets. A task-wise breakdown of performance is provided in Supplementary Tables S4 and S5.
@@
Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$)
@@
Interestingly, models trained with RL, particularly HuatuoGPT-o1-8B (RL), show the least degradation in performance compared to both standard SFT models and their own SFT-only variant (HuatuoGPT-o1-8B (SFT)). These results suggest that RL contributes to greater robustness against misleading initial reasoning and enhances backtracking capabilities. We provide further evidence supporting this claim in a later subsection (Table 4).
@@
Another interesting observation is that, reasoning-heavy questions tend to suffer larger performance drops than knowledge-heavy ones, as evidenced by the higher relative percentage decreases across most models. This pattern further highlights a gap in current models’ reasoning capabilities: while they may perform well on fact-based queries, their ability to reason through uncertainty, reconsider initial impressions, and course-correct remains limited.
Not all adversarial traces impact performance equally. As shown in Supplementary Table S6, Trace 1 leads to the largest performance drop compared to Traces 2 and 3, likely due to the greater uncertainty it introduces. These results suggest that different prompts induce varying degrees of reasoning disruption. While we evaluate only three examples, our findings highlight the importance of systematically exploring a diverse range of adversarial traces to better understand and improve model robustness under misleading or uncertain conditions.
4.4 Training Ablations
Building on insights from our evaluation, we examine how different training strategies impact a model’s reasoning ability and robustness. Specifically, we compare three paradigms—SFT, RL, and a hybrid SFT + RL approach—using reasoning-heavy and adversarial examples. This setup allows us to isolate the contribution of each method toward improving reasoning accuracy, resilience to misleading prompts, and generalization across medical QA benchmarks.
Model Knowledge Reasoning Overall
Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$)
Llama w/ m23k SFT 60.5 49.9 (-17.5) 47.1 35.8 (-24.0) 56.1 45.3 (-19.3)
+ Random Hard SFT 60.0 46.6 (-22.4) 47.9 31.2 (-34.9) 56.0 41.5 (-25.9)
+ Knowledge Hard SFT 60.1 46.7 (-22.3) 47.1 30.4 (-35.5) 55.8 41.4 (-25.9)
+ Reason Hard SFT 60.5 47.2 (-22.1) 49.1 30.0 (-38.9) 56.8 41.5 (-26.8)
+ Random Hard RL 62.6 51.5 (-18.1) 48.9 36.5 (-25.6) 58.1 46.4 (-20.2)
+ Knowledge Hard RL 62.7 51.4 (-18.1) 48.7 36.4 (-25.7) 58.2 46.5 (-20.2)
+ Reason Hard RL 63.6 52.1 (-18.0) 49.0 37.0 (-24.6) 58.8 47.2 (-19.8)
m23k + Reason Hard Full RL 59.5 42.5 (-28.6) 43.6 26.5 (-39.2) 54.3 37.3 (-31.4)
Table 4: Performance of Llama-3.1-8B-Instruct variants fine-tuned on m23k, followed by SFT, RL, or SFT+RL on 7,627 reasoning-hard examples. Controls use knowledge-hard and random-hard subsets. Reasoning-hard supervision yields the largest gains, especially under SFT. RL narrows differences across subsets, and SFT+RL consistently outperforms all other variants.
Figure 4: Average token counts generated by models trained with SFT and SFT + RL, stratified by reasoning-heavy and knowledge-heavy questions across all evaluation benchmarks. Models trained with RL on top of SFT tend to produce more concise reasoning traces compared to SFT-only models, despite both being trained on distilled traces from DeepSeek-R1. As expected, all models generate longer outputs for reasoning-heavy questions, reflecting their higher difficulty.
SFT + RL Enhances both Knowledge and Reasoning. Building on our finding that Llama fine-tuned on the m23k dataset achieves strong performance, we next examined whether augmenting training with carefully curated reasoning examples could further improve results. We combined three commonly used medical QA datasets—MedQA, MedMCQA, and HeadQA—and filtered 7,627 “reasoning-hard” multiple-choice questions using the following criteria: (i) neither Llama-3.1-8B-Instruct nor Qwen2.5-7B-Instruct answered them correctly more than 50% of the time, (ii) our PubMedBERT-based classifier assigned them a reasoning probability above 0.75, and (iii) they did not overlap with the m23k dataset. For comparison, we also curated two control sets of 7,627 examples each: a “knowledge-hard” set (with the same difficulty threshold but reasoning probability below 0.75) and a “random-hard” set (randomly sampled from similarly difficult examples). We generated detailed reasoning traces for all examples using R1 and trained models under three configurations: mixing m23k with reasoning-hard (+ Reason Hard SFT), with knowledge-hard (+ Knowledge Hard SFT), and with random-hard (+ Random Hard SFT). As shown in Table 4, incorporating reasoning-hard examples led to a measurable gain in reasoning accuracy (47.1% → 49.1%), while the knowledge-hard and random-hard sets yielded negligible improvements.
Finally, we applied the 7,627 reasoning-hard examples in an RL setup using GRPO (“+ Reason Hard RL”), continuing training from the m23k SFT checkpoint. For comparison, we also trained models using the random-hard and knowledge-hard subsets under the same RL setup (“+ Random Hard RL” and “+ Knowledge Hard RL”). RL training was conducted using the EasyR1 framework [47, 34], with key hyperparameters provided in Supplementary Table S1. Following RL, the performance gap between the ablation sets narrowed: all three variants showed improvements over their SFT-only counterparts, particularly in adversarial robustness. The “+ Reason Hard RL” model continued to perform best overall, achieving the highest scores across both knowledge and reasoning categories, under both normal and adversarial conditions. As an additional ablation, we also trained a model entirely with RL on the combined "m23k + Reason Hard" dataset (i.e., without any SFT), which yielded the worst performance—highlighting the importance of SFT initialization. Similar trends were observed with the Qwen2.5-7B-Instruct model, as shown in Supplementary Table S7. A detailed benchmark-wise breakdown is available in Supplementary Tables S8 and S9.
These results yield several important insights. While SFT on reasoning-hard examples primarily improves reasoning accuracy, RL produces broader gains across both reasoning- and knowledge-heavy tasks—likely because it encourages the model to prune irrelevant or inaccurate reasoning paths. Combining SFT with RL consistently outperforms either method alone. Although we do not claim to have identified the optimal balance, our findings suggest that some degree of SFT is especially important for general-domain models, which may lack sufficient exposure to biomedical content compared to domains like math or code. This observation aligns with prior work showing that SFT + RL typically yields the strongest results [15, 44, 10], though some studies have reported strong performance using RL-only training [6], possibly due to differences in model initialization or domain-specific pretraining.
While all models show performance degradation on out-of-distribution (OOD) benchmarks, those trained with SFT + RL exhibit stronger generalization—consistent with findings in Chu et al. [8]. These models also maintain higher accuracy across both knowledge and reasoning tasks, as shown in Supplementary Table S10. However, this effect appears to be model-dependent; for instance, Qwen2.5-7B-Instruct demonstrates a weaker improvement trend under the same conditions (Supplementary Table S11). A more systematic investigation into the role of model architecture and training dynamics is warranted, which we leave to future work.
Finally, we observe a significant difference in the number of tokens generated for reasoning-heavy versus knowledge-heavy problems. This is expected—models generally allocate more "thinking" tokens to harder tasks, and reasoning-heavy problems are empirically more difficult. Interestingly, models trained with an additional RL step tend to generate fewer tokens overall compared to their SFT-only counterparts, for both reasoning and knowledge-heavy queries, as shown in Supplementary Figure S6. This suggests that RL (specifically GRPO) may help models reason more efficiently, possibly by encouraging them to avoid verbose or redundant explanations while preserving correctness.
Model Knowledge Reasoning Overall
Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$)
Llama w/ m23k SFT 60.5 49.9 (-17.5) 47.1 35.8 (-24.0) 56.1 45.3 (-19.3)
Llama w/ m23k + Adv. SFT 60.8 58.0 (-4.6) 48.2 45.4 (-5.7) 56.6 53.8 (-4.9)
Qwen w/ m23k SFT 55.9 45.2 (-19.1) 41.3 31.7 (-23.3) 51.1 40.8 (-20.2)
Qwen w/ m23k + Adv. SFT 57.3 54.4 (-5.1) 42.8 40.8 (-4.8) 52.6 49.9 (-5.0)
Table 5: Training ablation including adversarial reasoning traces during SFT. Introducing a small proportion of such traces ($\sim$25%) significantly improves model robustness under uncertainty, reducing performance degradation in adversarial settings.
Including Adversarial Reasoning Traces Enhances Performance and Improves Robustness. To improve model robustness in adversarial settings, we explored whether training with uncertainty-inducing adversarial reasoning traces could help models recover from incorrect initial assumptions. We constructed a diverse set of adversarial trace templates that mimic varying degrees of uncertainty and hesitation—ranging from mild second-guessing to strong reliance on external but unreliable sources [49]. These prefills, shown in Supplementary Figure S15, include plausible but incorrect hypotheses (e.g., “I think the answer is {wrong answer}, but I’m not sure. Let me think again.”) designed to challenge the model’s reasoning trajectory. During training, each example in the m23k dataset had a 25% probability of being augmented with one of these adversarial traces, randomly sampled from our pool (Supplementary Figure S15). The selected prefill was appended to the beginning of the distilled reasoning trace, thereby teaching the model to override misleading signals and reason through uncertainty. As shown in Table 5, this simple augmentation strategy led to substantial gains in robustness. For instance, Llama’s performance drop under adversarial prompts decreased from 19.3% to just 4.9%, and Qwen’s dropped from 20.2% to 5.0%. The improvement was not limited to adversarial settings—standard accuracy also improved slightly across all metrics. This suggests that exposure to uncertain or misleading cues during training may help models better validate and correct their reasoning.
Model Knowledge Reasoning Overall
Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$) Normal Adv. (%$\Delta$)
m1-7B-23K 58.5 46.5 (-20.6) 45.5 34.2 (-24.7) 54.3 42.5 (-21.7)
HuatuoGPT-o1-8B 56.9 50.1 (-12.0) 44.8 35.3 (-21.2) 52.9 45.2 (-14.6)
OpenBioLLM-8B 45.4 16.8 (-63.0) 31.5 11.1 (-64.6) 40.8 14.9 (-63.4)
UltraMedical-8B 51.1 29.6 (-42.0) 39.2 18.9 (-51.8) 47.2 26.1 (-44.7)
MedReason-8B 58.2 28.2 (-51.5) 44.8 16.5 (-63.1) 53.8 24.4 (-54.7)
BioMed-R1-8B (Ours) 63.3 60.6 (-4.2) 49.2 47.7 (-2.9) 58.6 56.4 (-3.8)
MedGemma-27b 67.0 58.4 (-12.8) 53.9 45.2 (-16.2) 62.6 54.0 (-13.7)
m1-32B-1K 65.5 55.3 (-15.6) 53.7 42.3 (-21.3) 61.6 51.0 (-17.2)
HuatuoGPT-o1-70B 70.9 65.2 (-8.1) 56.9 50.0 (-12.1) 66.3 60.2 (-9.2)
OpenBioLLM-70B 63.1 47.3 (-25.0) 46.6 28.4 (-39.2) 57.7 41.1 (-28.8)
UltraMedical-70B 67.8 57.1 (-15.8) 52.4 41.2 (-21.3) 62.9 51.9 (-17.5)
BioMed-R1-32B (Ours) 68.2 65.4 (-4.1) 56.5 52.8 (-6.6) 64.4 61.3 (-4.8)
Table 6: Comparison of open-source biomedical reasoning models with our variants: BioMed-R1-8B and BioMed-R1-32B. BioMed-R1-8B leads among similarly sized models across all settings. BioMed-R1-32B rivals the much larger HuatuoGPT-o1-70B in standard accuracy and achieves the best adversarial robustness overall. All models are trained from public HuggingFace checkpoints.
Final Model Comparison and Takeaways. Building on insights from earlier experiments—namely, that RL consistently improves performance over SFT alone, reasoning-hard examples offer stronger supervision than other variants, and adversarial fine-tuning enhances robustness—we train our final models: BioMed-R1-8B and BioMed-R1-32B. BioMed-R1-8B is based on Llama-3.1-8B-Instruct, while BioMed-R1-32B is based on Qwen2.5-32B-Instruct. Both models are initialized with SFT on the m23k dataset, augmented with adversarial traces on 25% of examples, and further trained with RL on 7.5K reasoning-heavy questions filtered from MedQA, MedMCQA, and HeadQA.
Table 6 presents a comprehensive comparison against widely-used open-source biomedical reasoning models across knowledge-heavy, reasoning-heavy, and overall subsets—under both standard and adversarial prompting conditions. BioMed-R1-8B achieves the strongest performance among models of similar scale across all settings, highlighting the benefits of targeted RL and adversarial-aware training. BioMed-R1-32B ranks second in overall accuracy under standard prompts—closely matching HuatuoGPT-o1-70B despite being trained on substantially less data—and delivers the best adversarial robustness across all model sizes. As shown in Supplementary Figure S6, both models strike a strong balance in generation length: they avoid excessive verbosity while retaining sufficient reasoning detail. However, token counts for knowledge-based problems remain somewhat inflated. Future work could explore adaptive token budgeting strategies based on problem type to further improve efficiency. A detailed task-level breakdown is provided in Supplementary Table S12.